Government survey takers determine that typical family expenditures each month in the year designated as the base year are as follows:

• 20 pizzas, $10 each • Rent of apartment, $600 per month

• Gasoline and car maintenance, $100 per month

• Phone service (basic service plus 10 long-distance calls), $50 per month

In the year following the base year, the survey takers determine that pizzas have risen to $11 each, apartment rent is $640, gasoline and maintenance have risen to $120, and phone service has dropped in price to $40. Instructions: Enter your responses by rounding the CPI to three decimal places and the rate of inflation to one decimal place.

a. Find the CPI in the subsequent year and the rate of inflation between the base year and the subsequent year.

CPI:
Answer #1

Part A

Cost of the goods (base year) = ($20 × 10) + $600 + $100 + $50 = $950.

Cost of goods sold (Subsequent year) = ($20 × 11) + $640 + $120 + $40 = $1,020.

CPI in subsequent year = cost of the basket in that year relative to the base year = 1020/950 = 1.074

Rate of inflation = % increase in the CPI between the base year and the subsequent year = 1.074 – 1.000 = 0.074 = 7.4%